OpenStudy (anonymous):

how fast is the area of a square increasing at the instant when the sides are 4ft long if the sides increased at a rate of 1.5 ft/min

OpenStudy (anonymous):

\[a = s^2\]\[\frac{da}{dt} = 2s \frac{ds}{dt}\]

OpenStudy (anonymous):

thanks